### v1

# Low energy phonons in single crystal ZrW2O8

Abstract
ZrW2O8 is the prototypical example of a material exhibiting negative thermal expansion (NTE). As well as being among the first NTE materials to be studied in detail, it has served as a test bed for ideas for the physical mechanisms that lead to this unusual behavior. It is now widely accepted that here, and in many other framework materials exhibiting NTE, a collection of low energy phonon modes, as opposed to just one or two, are responsible for the anomalous thermal properties. However, quantitative verification and analysis of the density functional theory (DFT) calculations which underpin this proposal are still lacking. In particular, probing the low energy phonons directly throughout reciprocal space using inelastic neutron scattering, as opposed to other techniques which only probe the Brillouin zone center, is technically challenging and hence rarely done. Here we report inelastic neutron scattering measurements in a large number of Brillouin zones, achieved via the time of flight technique, over a 400 K temperature range. We find excellent agreement between DFT calculations and experimental data at low temperature, albeit with a rescaling of the calculated phonon energies, and of the lattice parameter used in the calculations compared to the measured value. However, the shifts in phonon modes due to the reduction in lattice parameter on warming, that can be predicted using DFT, are not observed. This is most likely due to counteractive anharmonic effects arising from the increased amplitude of lattice vibrations at elevated temperatures. Notwithstanding, the good agreement between experiment and DFT gives high confidence that the theory’s predictions of atomic motions for different modes is correct, and hence can be used as a reliable basis for quantitative analysis of NTE effects.
I Introduction
Negative thermal expansion (NTE) is when a material contracts rather than expands on warming. This effect can be uniform, i.e. along all crystallographic axes [1, 2, 3], or non-uniform [4, 5, 6]. It may also exist over a wide or a more limited range of temperatures [7]. ZrW2O8 is in many respects the canonical example, with its NTE having been re-discovered in the 1990s [8, 9, 10]. It is particularly notable that its NTE behavior is uniform, extant over a wide temperature range, from $\sim 10$ K to $\sim 1000$ K, and its coefficient of thermal expansion is large and negative ($\sim 10^{-5}\rm{K}^{-1}$) (see review articles ref. 11 and ref. 12 and references therein). In common with many other materials known to exhibit NTE, ZrW2O8 has a framework structure, i.e. it contains many ions that form coordinated polyhedra, most of which share corners, with large pores or spaces between the polyhedra. Indeed, it has more recently been found that intercalation of water into these pores can also result in significant anomalous lattice expansion or contraction [13]. The low temperature crystal structure of ZrW2O8 is shown in figure 1.
Figure 1: Crystal structure of ZrW2O8 with ZrO6 octahedra shaded green and WO4 tetrahedra shaded purple and blue, corresponding to tetrahedra centred on the crystallographically distinct W1 and W2 sites respectively. Note that each WO4 tetrahedron contains one unbonded oxygen ion on its vertex, pointing along a local $\langle$1 1 1$\rangle$-direction. The Zr–O–W2 bond angle is $\sim 172^{\circ}$ while the Zr–O–W1 bond angle is $\sim 154^{\circ}$. This figure was created using VESTA [14].
As a material is warmed phonon modes with gradually increasing energy are populated. If the potentials for these modes are asymmetric as a function of displacement, then thermal expansion will arise [15, 16]. Studies of the phonons in NTE materials often involve measurement of the phonon density of states (PDOS) using either Raman spectroscopy, inelastic x-ray scattering, or inelastic neutron scattering measurements of powder samples [17, 18, 19, 20, 21, 22, 23], as this is relatively straightforward to do and can be compared directly to harmonic lattice dynamics calculations such as those from density functional theory (DFT). ZrW2O8 is no exception, with the phonon density of states having been measured as a function of temperature soon after the seminal works on the NTE behavior [24]. It was found that the low energy peaks in the PDOS shift to higher energy on warming. Qualitatively, this is the opposite behavior to normal (positive thermal expansion) in which increased bond lengths on warming result in weaker interatomic potentials and hence a reduction in the energy of phonon modes.
In terms of interpretation of the PDOS measurements, and NTE more generally, early theories concerned the role of rigid unit modes (RUMs), in which Zr–O and W–O polyhedra remain undistorted but the whole structure flexes considerably along their shared vertices. As the amplitude of such modes increases, the root-mean-square distance between metal centers decreases, even though individual metal–oxygen bonds do not change in length [11]. Such modes would typically be found at the lower energy end of the phonon spectrum. However, various measurements have indicated that the picture is somewhat more complex than this. In particular, it has been found from EXAFS [25, 26, 27] that all of the polyhedra in the structure distort somewhat on warming, although the WO4 tetrahedra are noticeably more rigid than the ZrO6 octahedra. Notwithstanding, significant flex occurs along some of the Zr–O–W bonds that comprise the corner-sharing polyhedra. In this context, it would be appealing to search for a particular phonon mode that facilitates this distortion. However, numerous computational studies of the lattice dynamics in ZrW2O8 [28, 29, 30, 31] have demonstrated that there is no such single mode responsible for NTE. Rather, a large number of low energy phonon modes distributed throughout the Brillouin zone have a character that leads to NTE.
The direct measurements of phonons in ZrW2O8 (and many other NTE materials) to date have had a number of limitations. Raman spectroscopy can only measure the $\mathbf{q}=0$ Raman-active modes and thus does not provide a comprehensive picture, although the energy resolution available with this technique is usually superior to that available with neutron or x-ray spectroscopy. Neutron (or x-ray) scattering PDOS measurements are by their very nature a spherical average over reciprocal space of all possible phonon modes at a particular energy, so there is no direct sensitivity to the Q-space origin of shifts in energy with temperature. This can be ameliorated somewhat by combining and interpreting these measurements with DFT calculations. However, this approach is challenging in systems containing many atoms, and hence many phonon modes which may be more closely spaced than the energy resolution of the neutron spectrometer used for the measurements. In contrast, a measurement of a single crystal sample would allow individual modes to be isolated via measurements performed in many different Brillouin zones [32]. This would then provide a more stringent test of the DFT.
The motivation of the present study is therefore to develop a more comprehensive understanding of the low energy phonons in ZrW2O8 , by measuring in more Brillouin zones and with good energy resolution, to identify which particular individual or groups of phonon modes that are especially sensitive to temperature. Such measurements are sensitive to both the eigenvalues and eigenvectors extracted from DFT calculations, and hence pose a stringent test of the range of their validity, most notably for the low energy phonons thought to be important for NTE effects.
II Methods
ZrW2O8 crystallizes in the $P2_{1}3$ cubic space group, with a lattice parameter at room temperature of 9.16 Å[9]. The single-crystalline samples for these experiments were grown as described previously [33]. The sample used was composed of one crystal of mass 2.4 g with a mosaic of $\sim 0.3$ degrees, which is substantially less than the divergence of the incident neutron beam in the inelastic neutron scattering experiments described below, and hence unimportant for the experimental resolution. The lattice parameter was measured to be 9.176 Å  at 2 K, 9.159 Å  at 100 K, 9.1389 Å  at 300 K, and 9.125 Å  at 400 K during the alignment procedure for the inelastic scattering measurements described below. These values are consistent with those reported in the literature for powder samples [9].
II.1 Density functional theory
DFT calculations were performed using the plane-wave pseudopotential method as implemented in the CASTEP package [34]. The Perdew-Burke-Ernzehof (PBE) exchange-correlation functional [35] was used with norm-conserving pseudopotentials of the RRKJ type [36] from the Rappe and Bennett library [37], a plane-wave cutoff energy of 750 eV, and a $3\times 3\times 3$ Monkhorst-Pack [38] grid for the Brilloun-Zone of integration of electronic states. As also observed by Rimmer et al [28], the equilibrium lattice parameter at zero pressure is 9.259 Å, slightly larger than the measured low temperature lattice constant of $\sim 9.18$, typical of the PBE functional [39].
The phonon dispersion was calculated using Density Functional Perturbation Theory (DFPT) [40], with a $3\times 3\times 3$ Monkhorst-Pack grid of phonon wavevectors and Fourier interpolation of the resulting dynamical matrices, in order to sample the entire Brillouin zone. Additional calculations were performed for fixed unit cells of smaller dimension (9.23, 9.22, 9.215, 9.21 and 9.16 Å), at each of which the structure was relaxed and phonons calculated in the same way as for the fully relaxed structure. The calculated stress for each of these cell dimensions was 0.81, 1.13, 1.29, 1.45 and 3.20  GPa respectively. In the first four of these cases no soft phonon modes were observed, indicating that the computed structure remained dynamically stable. For the calculation with a lattice parameter of 9.16 Å a large number of negative energy modes were observed, indicating that for such a compressed lattice the structure becomes unstable. From this point onwards when we refer to these DFT calculations we mean those with $9.21\leq a\leq 9.23$.
When comparing the computed dispersion with that measured experimentally, an empirical scale factor is applied to the eigenvalues (energies) but not the eigenvectors (structure factors). This is determined by finding the simulation that agrees most closely with the data after such a uniform re-scale: the structure factors from the simulations with different lattice parameters are fitted to the neutron data with a single adjustable parameter (the energy scale). For ZrW2O8 the best overall agreement was found for the case of $a=9.215$and a scale factor $\sim 0.95$. This approach is used because changing the lattice parameter for the simulation does not affect the eigenvalues uniformly.
Additional calculations were performed with the same DFT parameters and a lattice parameter of 9.215 Å, but using the PBESol functional [39], which has been reported as yielding values for the bulk properties (elastic constants) of ZrW2O8  closer to the experimentally determined values than other functionals [41].
II.2 Inelastic neutron scattering
The time of flight (ToF) neutron scattering experiments were performed on the IN5 spectrometer at the Institut Laue-Langevin [42]. Data were collected with incident neutron energy $E_{i}=8.24$ meV, with the final chopper spun at 16,000 r.p.m. to give a resolution at the elastic line of 0.27 meV (full-width half-maximum, FWHM). The sample was aligned so that the $(H,H,0)$ and $(0,0,L)$ reciprocal space directions were in the horizontal plane of the instrument. The data were collected by scanning the sample orientation about the vertical axis in 0.5 degree steps over a range of 161 degrees, starting from the $(1,1,0)$-axis perpendicular to the incident neutron beam. Data were collected at temperatures of 100 K and 400 K, with these temperatures chosen in order to increase the signal via the Bose-Einstein population factor while maintaining a large temperature difference between measurements, but avoiding the structural phase transition at $\sim 425$ K [9]. Additional data were collected with the sample aligned with the $(H,0,0)$ and $(0,K,0)$ directions in the horizontal plane of the spectrometer, with the sample orientation scanned over a range of 88 degrees in 1 degree steps. These data were collected with $E_{i}=10.43$ meV and a final chopper speed of 16,000 r.p.m. giving a FWHM energy resolution at the elastic line of 0.35 meV. Data with this configuration were collected at temperatures of 2 K and 295 K.
The raw data were processed (units conversion, normalisation of intensities, etc.) using the Mantid software package [43]. Further processing of the data, and all subsequent visualisation and analysis, including convolution of calculated spectra with the instrumental resolution, were performed using the Horace software package [44]. Calculation of the inelastic neutron scattering cross-section ($S(\mathbf{Q},\omega)$) from the DFT-calculated phonons was done using the Euphonic software package’s Horace plugin [45, 46].
III Results
III.1 INS data
Figure 2 shows slices through the 4-dimensional datasets recording $S(\mathbf{Q},\omega)$, in the $\mathbf{Q}$–Energy plane, following a path through reciprocal space traversing a set of high-symmetry points. Note that although the points labelled as X, Y, and Z; and as M1, M2 and M3 are equivalent by symmetry, there are paths between them that are inequivalent by symmetry because the space group, $P2_{1}3$, is non-centrosymmetric. Data are presented centred at $\mathbf{G}=(0,0,4)$ in fig. 2 and $\mathbf{G}=(-2,-2,2)$ in fig. 3 for both $T=100$ K and $T=400$ K, and at $\mathbf{G}=(3,-3,0)$ in fig. 4 for $T=2$ K and $T=295$ K. The DFT-calculated dispersion relation with $a=9.215$, with the energy scale reduced by 5%, is overplotted (see discussion in section II.1 for further details). The calculated structure factor is also shown in panel (c) of all three figures. This set of figures illustrate many of the salient points of this work. It is clear that the dispersion and cross-section calculated using DFT provide an extremely good level of agreement with the neutron measurements, with significant overlap at most wavevectors, indicating that both the eigenvalues and eigenvectors from the DFT are accurate.
[…]
Figure 3: Slices through the $\mathbf{Q}$–Energy plane, following the same scheme as fig. 2. The data here were centered on the $\mathbf{Q}=(-2,-2,2)$ Brillouin zone. Notice that the change in direction of $\mathbf{Q}$ results in the highest intensity being peaked at different high symmetry positions than for the data shown in fig. 2.
Figure 4: Slices through the $\mathbf{Q}$–Energy plane, following the same scheme as fig. 2, with the exception of the extra reciprocal lattice point on the right hand side that is labelled M3 and refers to ($\frac{1}{2},-\frac{1}{2},0)$. The data here were collected in a separate experiment and are centered on the $\mathbf{Q}=(3,-3,0)$ Brillouin zone. Data shown were collected at 2 K and 300 K, which is different to figs. 2 and 3. The different experiments also have a different (but mutually consistent) set of arbitrary units for the intensity scale. These data are notable by having intensity peaked near the G point, unlike the other two datasets shown, and extending to slightly higher energy.
At most wavevectors a change in temperature of $\sim 300$ K results in very little change in the measured dispersion and $S(\mathbf{Q},\omega)$. A broadening of the signal from most of the phonons modes, indicative of anharmonicity at elevated temperatures, is observed. One notable exception is that on warming from 2 K to 300 K a significant mode hardening of around 0.5 meV is seen on the mode between the Y and R points in fig. 4. To show this in more detail a cut at the mid-point, $\mathbf{Q}=(3.25,-2.5,-0.25)$, is shown in fig. 5. Modes along similar trajectories in the reduced Brillouin zone in the other datasets (figs. 2 and 3) do not show such a dramatic shift. However, we can see from careful inspection of the DFT-calculated dispersion at the Y point that what appears to be a single mode is actually two different ones with only a small separation in energy. The fact that the two modes appear to behave differently with temperature is a good illustration of the power of measuring in multiple Brillouin zones - with only the data centered on $(-2,-2,2)$ and $(0,0,4)$ this effect would not have been visible. A similar, although less dramatic, mode hardening occurs at the mid-point of the Z–M2 trajectory (equivalent by symmetry to Y–M1), which is visible in both fig. 2 and 3, and is shown as a cut in fig. 6.
Figure 5: Cut at the mid-point of the Y–R trajectory of fig. 4(a) and (b), i.e. at $\mathbf{Q}=(3.25,-2.5,-0.25)$, showing a significant hardening of the first phonon peak from $\sim 3.5$ meV to $\sim 4$ meV.
Figure 6: Cut at the mid-point of the Z–M2 trajectory of fig. 2(a) and (b), i.e. at $\mathbf{Q}=(0,0.25,4.5)$, showing a slight hardening of the first phonon peak from $\sim 3.4$ meV to $\sim 3.5$ meV.
Despite the overall excellent agreement between the DFT calculations and the INS measurements shown in figs. 2, 3 and 4, there are a small number of wavevectors for which there is some disagreement. These can be seen by comparing the calculated dispersion and cross-section shown in panel (c) with the data in panels (a) and (b) of all three figures. For example, we can see that for the lowest energy mode on the path M1 – Y (and M2 – Z, which is equivalent by symmetry) the calculated energy is rather higher than the measurement at 100 K, whereas the other low energy ($<4$ meV) modes along other paths agree rather closely at this temperature.
To allow a more quantitative analysis we show constant $\mathbf{Q}$ cuts along the energy axis at various high symmetry wavevectors, G, X, M and R, at different temperatures in fig. 7. In all cases the integration range along the orthogonal $q$-axes is 0.04 reciprocal lattice units. This finite integration range, necessary when plotting data from time-of-flight INS experiments, means that for modes that are highly dispersive near the wavevector of interest the width of peaks may be broader than the nominal instrumental energy resolution. This is most apparent for the single peak in the cut at R – referring to fig. 3 one can see that the dispersion is very steep until quite close to the R point.
Figure 7: Cuts along the energy axis at four high symmetry positions in reciprocal space, centered on different Brillouin zones in order to maximize the cross-section. Panel (a) shows a cut at G, panel (b) a cut at Z, panel (c) a cut at M2, and panel (d) a cut at R. Data at lower temperatures (2 K in panel (a), 100 K in panels (b)–(d)) are shown in blue whereas data at higher temperatures (300 K in panel (a), 400 K in panels (b)–(d)) are shown in red. For all cuts an estimated constant background has been subtracted, followed by correction for the Bose-Einstein population factor in order to allow easier direct comparison of data taken at different temperatures. Solid lines correspond to fits to Gaussians (the number of which is set by the number expected from the DFT calculations), to indicate the peak position.
NTE in ZrW2O8 is reflected in the phonon dispersion relation via the mode Grüneisen parameters, $\gamma_{k}$. For a phonon mode $k$ with frequency $\omega_{k}$ this is defined as
$\gamma_{k}=-\frac{V}{\omega_{k}}\frac{\partial\omega_{k}}{\partial V}$ (1)
where $V$ is the volume of the unit cell. It was by analysis of the mode Gruüneisen parameters obtained from DFT calculations [28, 29, 30, 31] that it was proposed that NTE in ZrW2O8 is dominated by a collection of low energy phonon modes of varying character. However, the way mode Grüneisen parameters are determined from DFT is rather different to what actually happens when warming and cooling a sample. In the DFT, the lattice dynamics are calculated at zero temperature in the quasi-harmonic approximation for a series of different fixed lattice parameters close to the ground state parameters determined from a minimization of the overall forces and energies. This therefore corresponds to the lattice contracting or expanding under pressure / stress, rather than warming / cooling. In reality, the warming will result in an increased amplitude of lattice vibrations, which is not accounted for in the quasi-harmonic approximation, with the consequence that anharmonic effects may become more significant at higher temperatures. This point has been discussed extensively by Mittal et al [30, 12], who have calculated this effect for a few individual modes at particular wavevectors in ZrW2O8 . Some attempt has been made to quantify the effect more broadly using Raman spectroscopy data [17, 18], which has indicated that as well as having large negative Grüneisen parameters, the lowest energy modes also have the largest anharmonicity.
There are two effects which govern the behavior of a phonon mode energy with temperature. On the one hand, in a conventional material, thermal expansion (on increasing temperature) gives rise to longer bonds between atoms which are weaker and hence give rise to lower energy phonon modes. On the other hand, as temperature increases the amplitudes of normal modes increase and intrinsic anharmonicity comes into play. This would then yield phonon modes that harden slightly with increasing temperature [16]. However, in practice in conventional materials this effect is often much smaller than the softening due to thermal expansion and so is masked. In an NTE material one might imagine that if the lattice expansion effect is opposite to normal, then both of the effects described above would add together and result in phonon mode hardening substantially with increasing temperature. Or even in a RUM scenario, in which the bonds between atoms to not change length and hence strength, the intrinsic anharmonicity would be left over to give mode hardening, albeit perhaps smaller in magnitude. Neutron measurements of the phonon density of states in ZrW2O8 [24] indeed indicate that most of the low energy modes in ZrW2O8 do indeed harden on warming. Examining fig. 7, that shows cuts along energy at several key wavevectors in different Brillouin zones, we can gain greater insight into this.
At the G-point (fig. 7(a)) at 100 K, we observe one optic mode at $5.10\pm 0.03$ meV with a FWHM of $0.41\pm 0.06$ meV. On warming to 400 K the mode is fitted to be centered at $5.12\pm 0.05$ meV, with a FWHM of $0.85\pm 0.1$ meV. Our calculations, and those of other groups, indicate that one might expect to see two modes here (see fig. 4, for example). However they are close in energy and dispersive so may not be possible to resolve and may be broadened by additional instrumental effects beyond the bare energy resolution (which is $\sim 0.23$ meV at this energy). Certainly, the peak is noticeably broadened, most likely due to thermal effects, at 400 K. The behavior of this peak contrasts with the expected softening due to the change in lattice alone of -0.2 meV (for both calculated modes), which Gupta et al [30] predict to be compensated such that the expected shifts are $-0.07$ meV and $-0.03$ meV respectively, which is significantly less than the energy resolution of our measurements  [^1] Note that the predictions of Gupta et al [30] refer to a change in temperature of 0 K to 300 K, i.e. the same difference in temperature as for our measurements. The offset is unlikely to be important, since NTE in ZrW2O8 is linear over this temperature range.. The Raman measurements of Oishi et al include this mode [18], which they find to be centered at 5.16 meV and largely independent of temperature, consistent with our finding.
At the Z-point (fig. 7(b), equivalent to the X- and Y-points) we observe three modes. At 100 K they are centered at energies of $3.50\pm 0.05$ meV, $4.43\pm 0.05$ meV and $5.28\pm 0.04$ meV. All three were found to have the same FWHM (within error) of $0.49\pm 0.05$ meV. At 400 K these peaks are centered at energies of $3.70\pm 0.10$ meV, $4.48\pm 0.09$ meV and $5.44\pm 0.13$ meV respectively. The FWHMs were found to be $0.53\pm 0.11$ meV. Gupta et al presented predictions for the first two of these modes, with a shift due to the change in lattice of -0.14 and -0.16 meV respectively, that is compensated by thermal effects to give expected shifts of 0.22 meV and 0.68 meV. The first of these predictions is consistent with our measurement, whereas the latter is not – the predicted effect of temperature (referred to as explicit anharmonicity) is larger than what is actually measured.
At the M-point (fig. 7(c), actually M2 in our notation) we observe three modes. At 100 K they are centered at energies of $3.77\pm 0.04$ meV, $4.35\pm 0.05$ meV and $5.60\pm 0.03$ meV. All three were found to have the same FWHM (within error) of $0.42\pm 0.04$ meV. At 400 K these peaks are centered at energies of $3.58\pm 0.07$ meV, $4.05\pm 0.06$ meV and $5.62\pm 0.04$ meV respectively. The FWHMs were found to be $0.34\pm 0.06$ meV. For the first two of these modes, Gupta et al predict shifts due to a change in lattice of -0.36 and -0.37 meV respectively, compensated by explicit anharmonicity to give overall predicted shifts of 0.06 and 0.18 meV. In contrast, our fits would seem to indicate that both modes soften slightly. However, inspecting fig. 3 we see that care should be taken here – it is not obvious from the colormap that these two modes soften as much as the cuts and fits seem to suggest. In fact, what appears to happen is the lowest energy mode which disperses weakly along the Z – M2 trajectory with an energy of $\sim 3.5$ meV seems to steepen slightly on warming, with the effect of bringing some of its spectral weight into the integration region of the cut at 400 K shown in fig. 7(c). Therefore, in the absence of a calculation for the shifts in energy at all $\mathbf{Q}$, we cannot quantify with much accuracy how well the measurement and prediction agree at this wavevector. The predictions suggest small shifts at about the level of our experimental resolution, which looks qualitatively reasonable.
At the R-point (fig. 7(d)) we observe one mode. At 100 K it is centered at an energy of $4.45\pm 0.02$ meV, with a FWHM of $0.56\pm 0.03$ meV. At 400 K the peak is centered at an energy of $4.43\pm 0.05$ meV with a FWHM of $0.91\pm 0.11$ meV. Here, Gupta et al predict a shift due to the change in lattice of -0.39 meV, which is complemented by a significant explicit anharmonicity of the opposite sign to those at other wavevectors, to give an overall predicted shift of -0.77 meV. The large FWHM observed at both temperatures is due to the steep dispersion of the modes near the R-point and the consequent inclusion in the cut integration region of spectral weight away from the R-point itself. Notwithstanding this, it is clear from both fig. 7(d) and fig. 3 that the mode near R does not soften as much as predicted.
III.2 DFT calculations
The DFT calculations were performed at a lattice parameter of 9.215Å, and we applied a global re-scaling of the calculated phonon energies of 95% to give best agreement with the INS measurements  [^2] This was done by doing a least squares fit of the neutron spectra with the DFT-calculated spectra, varying only the scale factor to give the best agreement. Coincidentally, the relative difference in lattice parameter between that used for the DFT and that observed experimentally at low temperature is approximately the same as the experimentally observed negative thermal expansion between 100 K and 400 K [9]. We show in fig. 8 the effect of the lattice parameter on the computed phonon spectra below 6 meV (the energy range covered in our INS experiments). As would be expected from the Grüneisen parameters reported in the literature, the low energy phonons are mostly found to harden with increasing lattice parameter. However, the scaling is not uniform, as can be seen in the inset which shows the trajectory along Z – M2 – Y. Here it is clear that the first optic mode (labelled 2 and 4 for the simulation with $a=9.259$  (the optimized lattice parameter in the DFT calculations) and $a=9.215$  respectively) is significantly more dispersive between Z and M2 for the larger lattice parameter. This also illustrates why we chose to compare the INS data with the calculations for the unit cell with $a=9.215$  rather than $a=9.259$. A global re-scaling of the latter results in a worse agreement with the measurement than the former.
Figure 8: Comparison of dispersion from DFT calculations with different lattice parameter. The reciprocal space path is the same a described in fig. 2. The solid red curve is for calculations with lattice parameter $a=9.259$(zero stress, optimized lattice parameter), and the dashed blue curve is for calculations with $a=9.215$. The inset shows a zoomed in section of the dispersion between 3 and 6 meV. The modes labelled 1, 2 and 5 correspond to calculations with $a=9.215$, and the modes labelled 3, 4 and 6 correspond to calculations with $a=9.259$. The symbols at high symmetry wavevectors correspond to the fitted peak positions for the lower temperature datasets shown in fig. 7, with diamonds for the G-point, stars for the Z-point, pentagons for the M2-point and a circle at the R-point.
One might naturally ask if further compression of the lattice in the calculations could yield yet better agreement. In fig. 9 we show calculations with $a=9.215$  and $9.210$. Here we see that certain modes, for example on the path Y – R – G, have softened considerably, such that an overall re-scale of energies would once again not yield good agreement with the INS data. What this seems to illustrate is that the calculations which agree best with the INS data are those which are on the edge of dynamical stability. Indeed, as noted in sec. II.1, when the lattice was compressed further in the DFT calculations, a significant number of soft modes appeared, consistent with this interpretation. With further compression negative energy modes appear, indicating an eventual change in the ground state structure.
Figure 9: Comparison of dispersion from DFT calculations with lattice compressed further. The reciprocal space path is the same a described in fig. 2. The solid cyan curve is for calculations with lattice parameter $a=9.210$, and the dashed blue curve is for calculations with $a=9.215$.
Figure 10: The phonon density of states (PDOS) as a function of energy computed using DFT. The total PDOS is shown by the black curve, and the element-wise contributions are shown in red (zirconium), green (tungsten) and blue (oxygen).
Fig. 10 shows the partial phonon density of states (PDOS) computed from DFT, broken down into the contributions from each of the constituent atomic species. The advantage of computing this property is that it is intrinsically averaged across all of reciprocal space and hence offers some insight into the combined effect of many different phonon modes to the material’s thermal properties. The disadvantage, as already noted, is that the averaging means that information about particular phonon modes is lost.
In the energy range probed by this experiment, the PDOS has approximately equal weight from tungsten and oxygen, but very little contribution from zirconium. This indicates that the zirconium ions are much more fixed in position, somewhat characteristic of rigid-unit-modes. Further decomposition of the PDOS down to the individual atom level (see supplementary information [32]) shows that the contribution from individual tungsten ions is larger than for individual oxygen ions by about a factor two to three at low energies, but the oxygen ions are more numerous so sum up to give a comparable contribution to the PDOS.
Figure 11: The phonon density of states (PDOS) weighted by the element-specific neutron scattering factors, as a function of energy computed using DFT. The total PDOS is shown by the black curve, and the element-wise contributions are shown in red (zirconium), green (tungsten) and blue (oxygen).
Figure 12: Slices through the $\mathbf{Q}$ energy plane, following the same scheme as fig. 3. The colormaps show the neutron scattering intensity calculated using DFT, decomposed into the contributions from each individual element. Panel (a) shows the contribution from zirconium, panel (b) the contribution from tungsten, and panel (c) the contribution from oxygen. Note that the intensity scale is the same as that in fig. 3 and is the same for all three panels.
Fig. 11 shows the PDOS with neutron weighting applied (see eq. 1 in the supplementary information [32]), which allows direct comparison with the neutron structure factor measured in an experiment, such as that in ref. 24. The scattering lengths of all three constituent elements are similar in magnitude, as are their Debye-Waller factors, so the largest effect on the neutron cross-section comes from the division of all the other terms by the atomic mass. This gives significantly more weight to the oxygen modes than to tungsten and zirconium, with the consequence that the neutron scattering signal is dominated by oxygen motion. We can use a similar procedure for the single crystal plots shown in figs. 24, i.e. of element-wise decomposition of the signal. Examples are shown in fig. 12, for the $(-2,2,2)$ Brillouin zone center, i.e. equivalent to fig. 3. As with the PDOS, most of the signal arises from oxygen motions, with contributions from tungsten and zirconium around an order of magnitude smaller. However, although the neutron measurements are not directly sensitive to the motion of the heavier metal atoms, the fact that the DFT gives good agreement for the oxygen means that we can be confident that its predictions for the zirconium and tungsten motion are accurate, since each of them are surrounded by a cage of oxygen.
IV Discussion
From the results and analysis presented so far, we can see that there is generally very good agreement between the phonon dispersion relations and structure factors computed from DFT, and our inelastic neutron scattering measurements. Of course, anharmonicity is essential to describe any kind of thermal expansion, negative or positive. We have already cited extensively the predictions made by Gupta et al [30], which were based on a phenomenological theory developed for understanding librational modes at $q=0$ in the $\alpha$ phase of solid nitrogen [49]. In this approach the phonon modes are decomposed into the displacement and rotation of molecular units (analogous to the RUM approach) and the computed potential rewritten as a polynomial to $4^{\rm{th}}$ order in terms of these coordinates. This is clearly an approximation, though it should be acknowledged that it does a reasonable job, estimating correctly that in ZrW2O8 at some wavevectors the effects of implicit and explicit anharmonicity seem to approximately cancel each other.
It would seem, then, that for the low energy excitations studied here there is a cancellation of two large effects on the energy of the modes. Both arise due to anharmonicity, and given the unconventional behavior of ZrW2O8 it is clear that anharmonicity has an important role in governing the material’s thermal properties. In order to cancel the mode softening expected from the mode Grüneisen parameters, large amplitude thermal motions are required. This would naturally arise in a structure that is especially flexible, with individual atoms subject either only to weak constraints or strongly constrained only in one or two directions.
Alternative methods exist in addition to DFT of computing the phonon spectra, that include anharmonicity. These include the temperature-dependent effective potential (TDEP) method [50, 51, 52] and the self-consistent phonon (SCPH) method [53, 54]. These would in principle allow us to investigate in detail the effect of temperature and anharmonicity on the phonon modes, and hence NTE. However, the calculations to do this for a system like ZrW2O8 would be prohibitively computationally expensive, and are thus outside the scope of the present work.
Notwithstanding the aforementioned shortcomings of the DFT computations, it should be acknowledged that they perform an admirable job of reproducing many of the details of the INS measurements, both in terms of the eigenvalues (mode energies) and eigenvectors (measured neutron scattering intensities). This then gives high confidence that they can be used to give both a qualitative and quantitative commentary on the nature of the atomic displacements involved in the low energy phonon modes measured here. This was discussed extensively already by Gupta et al [30], who correlated the translational and rotational motions of quasi-rigid ZrO6 octahedra and WO4 tetrahedra, together with distortions of those polyhedra, with reported distortions found from local structure probes [55]. We can also have good confidence that the DFT results can be used as a general interpretative tool, e.g. in the flexibility analysis of NTE undertaken by Rimmer et al [28].
Molecular dynamics simulations [29] suggest that the oxygens that bridge Zr–O–W bonds are subject to large transverse vibrations. The distortions we observe from our DFT computations indicate that some modes do indeed include such motions. Because of the relative rigidity of the WO4 tetrahedra (it is well-known that the W–O bonds are shorter than the Zr–O bonds, so the WO4 tetrahedra are more tightly bonded and stiffer than the ZrO6 octahedra [56]) a consequence of such vibrations is that the lone unbonded oxygen on each tetrahedron is also subject to large vibrations [25]. Indeed, displacements of the lone oxygen would have a lower energy cost due to weaker bonding. The flexibility analysis performed by Rimmer et al [28], using DFT results similar to those reported here, showed that the low energy phonon modes responsible for most of the NTE (those with negative Grüneisen parameters) had largely fixed Zr–O and W–O bond lengths. These results together suggest that the tension effect description of NTE offers an appropriate description of the behavior of ZrW2O8 : rigid Zr–O and W–O bonds together with a large transverse vibration of oxygens that bridge Zr and W lead naturally to a reduced distance between Zr and W and hence NTE. By validating the DFT that underpins this interpretation against INS data, we are able to give greater credence to it. We also note that in the high pressure orthorhombic phase of $\gamma$-ZrW2O8 [57] the lone oxygen becomes bonded to a tungsten ion, making its motion more constrained than in the ambient pressure cubic $\alpha$-ZrW2O8 phase, which is then manifest in a significantly smaller (negative) coefficient of thermal expansion in the high pressure polymorph.
V Conclusion
In conclusion, we present a comprehensive set of INS measurements of the phonons in ZrW2O8 , and compare them to predictions from DFT. We find that we have to compress the lattice compared to the equilibrium in the DFT in order to get agreement with the neutron data, but that after doing so the agreement is excellent. This gives confidence that the overall picture of atomic motions predicted by DFT closely matches the experimental situation, so that DFT can be confidently used as a basis for additional analysis of NTE. The calculated mode Grüneisen parameters suggest that most low energy phonon modes should soften on warming as the lattice shrinks, but instead most modes either remain unchanged or harden slightly, as well as broadening somewhat. This can be understood in terms of anharmonicity arising from an increased amplitude of atomic motions associated with the phonon modes.
Acknowledgements.
We are grateful to D. Voneshen (ISIS) and R. Fair (STFC-SCD) for their advice and assistance on the use of Euphonic. We are indebted to M. Dove (QMUL) for his insightful and informative comments. Raw data are available via ref. [58]. We thank the Institut Laue-Langevin for the allocation of beamtime for the experiments reported here. Computing resources were provided by the STFC Scientific Computing Department’s SCARF cluster.

### v2

@@
 ZrW2O8 is the prototypical example of a material exhibiting negative thermal expansion (NTE).
-As well as being among the first NTE materials to be studied in detail, it has served as a test bed for ideas for the physical mechanisms that lead to this unusual behavior.
-It is now widely accepted that here, and in many other framework materials exhibiting NTE, a collection of low energy phonon modes, as opposed to just one or two, are responsible for the anomalous thermal properties.
+It is now widely accepted that in ZrW2O8, and in many other framework materials exhibiting NTE, a collection of low energy phonon modes, as opposed to just one or two, are responsible for the anomalous thermal properties.
 However, quantitative verification and analysis of the density functional theory (DFT) calculations which underpin this proposal are still lacking.
 In particular, probing the low energy phonons directly throughout reciprocal space using inelastic neutron scattering, as opposed to other techniques which only probe the Brillouin zone center, is technically challenging and hence rarely done.
-Here we report inelastic neutron scattering measurements in a large number of Brillouin zones, achieved via the time of flight technique, over a 400 K temperature range.
-We find excellent agreement between DFT calculations and experimental data at low temperature, albeit with a rescaling of the calculated phonon energies, and of the lattice parameter used in the calculations compared to the measured value.
-However, the shifts in phonon modes due to the reduction in lattice parameter on warming, that can be predicted using DFT, are not observed.
-This is most likely due to counteractive anharmonic effects arising from the increased amplitude of lattice vibrations at elevated temperatures.
-Notwithstanding, the good agreement between experiment and DFT gives high confidence that the theory’s predictions of atomic motions for different modes is correct, and hence can be used as a reliable basis for quantitative analysis of NTE effects.
+Here we report inelastic neutron scattering measurements in a large number of Brillouin zones over a 400 K temperature range.
+We find excellent agreement between DFT calculations and experimental data at low temperature.
+However, the shifts in phonon modes predicted by DFT due to the reduction in lattice parameter (warming) are not observed.
+This is most likely due to counteractive anharmonic effects, which we verified using finite temperature molecular dynamics (MD) calculations.
+Notwithstanding, both DFT and MD results are consistent with NTE in ZrW2O8 arising from the tension effect, and by extension this explanation is supported by the neutron scattering results.
 I Introduction
@@
 In common with many other materials known to exhibit NTE, ZrW2O8 has a framework structure, i.e.
-it contains many ions that form coordinated polyhedra, most of which share corners, with large pores or spaces between the polyhedra.
-Indeed, it has more recently been found that intercalation of water into these pores can also result in significant anomalous lattice expansion or contraction [13].
-The low temperature crystal structure of ZrW2O8 is shown in figure 1.
-Figure 1: Crystal structure of ZrW2O8 with ZrO6 octahedra shaded green and WO4 tetrahedra shaded purple and blue, corresponding to tetrahedra centred on the crystallographically distinct W1 and W2 sites respectively.
+it contains many ions that form coordinated polyhedra, most of which share corners, with large pores or spaces between the polyhedra (see fig.
+1.
+It has also been found that intercalation of water into these pores can also result in significant anomalous lattice expansion or contraction [13].
+Figure 1: The crystal structure of ZrW2O8 (space group $P2_{1}3$) with ZrO6 octahedra shaded green and WO4 tetrahedra shaded purple and blue, corresponding to tetrahedra centred on the crystallographically distinct W1 and W2 sites respectively.
 Note that each WO4 tetrahedron contains one unbonded oxygen ion on its vertex, pointing along a local $\langle$1 1 1$\rangle$-direction.
-The Zr–O–W2 bond angle is $\sim 172^{\circ}$ while the Zr–O–W1 bond angle is $\sim 154^{\circ}$.
 This figure was created using VESTA [14].
-As a material is warmed phonon modes with gradually increasing energy are populated.
+As a material is warmed, phonon modes with gradually increasing energy are populated.
 If the potentials for these modes are asymmetric as a function of displacement, then thermal expansion will arise [15, 16].
-Studies of the phonons in NTE materials often involve measurement of the phonon density of states (PDOS) using either Raman spectroscopy, inelastic x-ray scattering, or inelastic neutron scattering measurements of powder samples [17, 18, 19, 20, 21, 22, 23], as this is relatively straightforward to do and can be compared directly to harmonic lattice dynamics calculations such as those from density functional theory (DFT).
+Studies of the phonons in NTE materials often involve measurement of the phonon density of states (PDOS) using either Raman spectroscopy, inelastic x-ray scattering, or inelastic neutron scattering measurements of powder samples [17, 18, 19, 20, 21, 22, 23], as this is relatively straightforward to do and can be compared directly to quasi-harmonic lattice dynamics calculations such as those from density functional theory (DFT).
 ZrW2O8 is no exception, with the phonon density of states having been measured as a function of temperature soon after the seminal works on the NTE behavior [24].
 It was found that the low energy peaks in the PDOS shift to higher energy on warming.
-Qualitatively, this is the opposite behavior to normal (positive thermal expansion) in which increased bond lengths on warming result in weaker interatomic potentials and hence a reduction in the energy of phonon modes.
+Qualitatively, this is the opposite behavior to normal (positive thermal expansion) in which increased bond lengths on warming typically result in weaker interatomic potentials and hence a reduction in the energy of phonon modes.
 In terms of interpretation of the PDOS measurements, and NTE more generally, early theories concerned the role of rigid unit modes (RUMs), in which Zr–O and W–O polyhedra remain undistorted but the whole structure flexes considerably along their shared vertices.
@@
 Such modes would typically be found at the lower energy end of the phonon spectrum.
-However, various measurements have indicated that the picture is somewhat more complex than this.
+However, various measurements have indicated that the picture may be somewhat more complex than this.
 In particular, it has been found from EXAFS [25, 26, 27] that all of the polyhedra in the structure distort somewhat on warming, although the WO4 tetrahedra are noticeably more rigid than the ZrO6 octahedra.
@@
 The direct measurements of phonons in ZrW2O8 (and many other NTE materials) to date have had a number of limitations.
-Raman spectroscopy can only measure the $\mathbf{q}=0$ Raman-active modes and thus does not provide a comprehensive picture, although the energy resolution available with this technique is usually superior to that available with neutron or x-ray spectroscopy.
+Raman spectroscopy can only measure the Brillouin zone center ($\mathbf{q}=0$) Raman-active modes and thus does not provide a comprehensive picture, although the energy resolution available with this technique is usually superior to that available with neutron or x-ray spectroscopy.
 Neutron (or x-ray) scattering PDOS measurements are by their very nature a spherical average over reciprocal space of all possible phonon modes at a particular energy, so there is no direct sensitivity to the Q-space origin of shifts in energy with temperature.
-This can be ameliorated somewhat by combining and interpreting these measurements with DFT calculations.
+This can be ameliorated somewhat by interpreting these measurements with DFT calculations.
 However, this approach is challenging in systems containing many atoms, and hence many phonon modes which may be more closely spaced than the energy resolution of the neutron spectrometer used for the measurements.
-In contrast, a measurement of a single crystal sample would allow individual modes to be isolated via measurements performed in many different Brillouin zones [32].
+In contrast, a measurement of a single crystal sample would allow individual modes to be isolated via measurements performed in many different Brillouin zones (see appendices for a more detailed discussion).
 This would then provide a more stringent test of the DFT.
-The motivation of the present study is therefore to develop a more comprehensive understanding of the low energy phonons in ZrW2O8 , by measuring in more Brillouin zones and with good energy resolution, to identify which particular individual or groups of phonon modes that are especially sensitive to temperature.
-Such measurements are sensitive to both the eigenvalues and eigenvectors extracted from DFT calculations, and hence pose a stringent test of the range of their validity, most notably for the low energy phonons thought to be important for NTE effects.
+The motivation of the present study is therefore to develop a more comprehensive understanding of the low energy phonons in ZrW2O8 , by measuring in more Brillouin zones using inelastic neutron scattering with good energy resolution, to identify if any particular individual or groups of phonon modes are especially sensitive to temperature.
+Such measurements are sensitive to both the eigenvalues and eigenvectors extracted from ab-initio calculations, and hence pose a stringent test of the range of their validity, most notably for the low energy phonons thought to be important for NTE.
 II Methods
-ZrW2O8 crystallizes in the $P2_{1}3$ cubic space group, with a lattice parameter at room temperature of 9.16 Å[9].
-The single-crystalline samples for these experiments were grown as described previously [33].
-The sample used was composed of one crystal of mass 2.4 g with a mosaic of $\sim 0.3$ degrees, which is substantially less than the divergence of the incident neutron beam in the inelastic neutron scattering experiments described below, and hence unimportant for the experimental resolution.
-The lattice parameter was measured to be 9.176 Å  at 2 K, 9.159 Å  at 100 K, 9.1389 Å  at 300 K, and 9.125 Å  at 400 K during the alignment procedure for the inelastic scattering measurements described below.
-These values are consistent with those reported in the literature for powder samples [9].
-II.1 Density functional theory
-DFT calculations were performed using the plane-wave pseudopotential method as implemented in the CASTEP package [34].
-The Perdew-Burke-Ernzehof (PBE) exchange-correlation functional [35] was used with norm-conserving pseudopotentials of the RRKJ type [36] from the Rappe and Bennett library [37], a plane-wave cutoff energy of 750 eV, and a $3\times 3\times 3$ Monkhorst-Pack [38] grid for the Brilloun-Zone of integration of electronic states.
-As also observed by Rimmer et al [28], the equilibrium lattice parameter at zero pressure is 9.259 Å, slightly larger than the measured low temperature lattice constant of $\sim 9.18$, typical of the PBE functional [39].
-The phonon dispersion was calculated using Density Functional Perturbation Theory (DFPT) [40], with a $3\times 3\times 3$ Monkhorst-Pack grid of phonon wavevectors and Fourier interpolation of the resulting dynamical matrices, in order to sample the entire Brillouin zone.
-Additional calculations were performed for fixed unit cells of smaller dimension (9.23, 9.22, 9.215, 9.21 and 9.16 Å), at each of which the structure was relaxed and phonons calculated in the same way as for the fully relaxed structure.
-The calculated stress for each of these cell dimensions was 0.81, 1.13, 1.29, 1.45 and 3.20  GPa respectively.
-In the first four of these cases no soft phonon modes were observed, indicating that the computed structure remained dynamically stable.
-For the calculation with a lattice parameter of 9.16 Å a large number of negative energy modes were observed, indicating that for such a compressed lattice the structure becomes unstable.
-From this point onwards when we refer to these DFT calculations we mean those with $9.21\leq a\leq 9.23$.
-When comparing the computed dispersion with that measured experimentally, an empirical scale factor is applied to the eigenvalues (energies) but not the eigenvectors (structure factors).
-This is determined by finding the simulation that agrees most closely with the data after such a uniform re-scale: the structure factors from the simulations with different lattice parameters are fitted to the neutron data with a single adjustable parameter (the energy scale).
-For ZrW2O8 the best overall agreement was found for the case of $a=9.215$and a scale factor $\sim 0.95$.
-This approach is used because changing the lattice parameter for the simulation does not affect the eigenvalues uniformly.
-Additional calculations were performed with the same DFT parameters and a lattice parameter of 9.215 Å, but using the PBESol functional [39], which has been reported as yielding values for the bulk properties (elastic constants) of ZrW2O8  closer to the experimentally determined values than other functionals [41].
-II.2 Inelastic neutron scattering
-The time of flight (ToF) neutron scattering experiments were performed on the IN5 spectrometer at the Institut Laue-Langevin [42].
-Data were collected with incident neutron energy $E_{i}=8.24$ meV, with the final chopper spun at 16,000 r.p.m.
-to give a resolution at the elastic line of 0.27 meV (full-width half-maximum, FWHM).
-The sample was aligned so that the $(H,H,0)$ and $(0,0,L)$ reciprocal space directions were in the horizontal plane of the instrument.
-The data were collected by scanning the sample orientation about the vertical axis in 0.5 degree steps over a range of 161 degrees, starting from the $(1,1,0)$-axis perpendicular to the incident neutron beam.
-Data were collected at temperatures of 100 K and 400 K, with these temperatures chosen in order to increase the signal via the Bose-Einstein population factor while maintaining a large temperature difference between measurements, but avoiding the structural phase transition at $\sim 425$ K [9].
-Additional data were collected with the sample aligned with the $(H,0,0)$ and $(0,K,0)$ directions in the horizontal plane of the spectrometer, with the sample orientation scanned over a range of 88 degrees in 1 degree steps.
-These data were collected with $E_{i}=10.43$ meV and a final chopper speed of 16,000 r.p.m.
-giving a FWHM energy resolution at the elastic line of 0.35 meV.
-Data with this configuration were collected at temperatures of 2 K and 295 K.
-The raw data were processed (units conversion, normalisation of intensities, etc.) using the Mantid software package [43].
-Further processing of the data, and all subsequent visualisation and analysis, including convolution of calculated spectra with the instrumental resolution, were performed using the Horace software package [44].
-Calculation of the inelastic neutron scattering cross-section ($S(\mathbf{Q},\omega)$) from the DFT-calculated phonons was done using the Euphonic software package’s Horace plugin [45, 46].
+Time of flight neutron scattering experiments were performed on the IN5 spectrometer at the Institut Laue-Langevin [32].
+Data were collected with incident neutron energies $E_{i}=8.24$ meV and $E_{i}=10.43$ meV, at temperatures in the range 2 K to 400 K.
+Density Functional Theory (DFT) calculations were performed using the plane-wave pseudopotential method as implemented in the CASTEP package [33].
+Calculations were performed for a series of different fixed lattice parameters close to the ground state parameters determined from a minimization of the overall forces and energies.
+Phonon dispersions were calculated using Density Functional Perturbation Theory (DFPT).
+It is important to note that in DFT the lattice dynamics are calculated at zero temperature in the quasi-harmonic approximation.
+When comparing the computed dispersion with that measured experimentally with neutrons, an empirical scale factor was applied to the eigenvalues (phonon energies) but not to the eigenvectors (structure factors).
+Calculations of the finite-temperature behavior of the phonon dispersion at a few high-symmetry wavevectors were performed using machine learned potentials and molecular dynamics (MD).
+The potentials were computed using the automated potential development (APD) workflow [34].
+These potentials were then used for MD simulations, using the LAMMPS software package [35].
+Further details of the inelastic neutron scattering experimental methods can be found in the appendices, as can the detailed schemes used for the density functional theory and molecular dynamics calculations.
 III Results
 III.1 INS data
-Figure 2 shows slices through the 4-dimensional datasets recording $S(\mathbf{Q},\omega)$, in the $\mathbf{Q}$–Energy plane, following a path through reciprocal space traversing a set of high-symmetry points.
-Note that although the points labelled as X, Y, and Z; and as M1, M2 and M3 are equivalent by symmetry, there are paths between them that are inequivalent by symmetry because the space group, $P2_{1}3$, is non-centrosymmetric.
-Data are presented centred at $\mathbf{G}=(0,0,4)$ in fig.
-2 and $\mathbf{G}=(-2,-2,2)$ in fig.
-3 for both $T=100$ K and $T=400$ K, and at $\mathbf{G}=(3,-3,0)$ in fig.
-4 for $T=2$ K and $T=295$ K.
-The DFT-calculated dispersion relation with $a=9.215$, with the energy scale reduced by 5%, is overplotted (see discussion in section II.1 for further details).
-The calculated structure factor is also shown in panel (c) of all three figures.
-This set of figures illustrate many of the salient points of this work.
-It is clear that the dispersion and cross-section calculated using DFT provide an extremely good level of agreement with the neutron measurements, with significant overlap at most wavevectors, indicating that both the eigenvalues and eigenvectors from the DFT are accurate.
 […]
@@
 Figure 4: Slices through the $\mathbf{Q}$–Energy plane, following the same scheme as fig.
-2, with the exception of the extra reciprocal lattice point on the right hand side that is labelled M3 and refers to ($\frac{1}{2},-\frac{1}{2},0)$.
+2, with the exception of the extra reciprocal lattice point on the right hand side that is labeled M3 and refers to ($\frac{1}{2},-\frac{1}{2},0)$.
 The data here were collected in a separate experiment and are centered on the $\mathbf{Q}=(3,-3,0)$ Brillouin zone.
@@
 These data are notable by having intensity peaked near the G point, unlike the other two datasets shown, and extending to slightly higher energy.
+Figure 2 shows slices through the 4-dimensional datasets recording $S(\mathbf{Q},\omega)$, in the $\mathbf{Q}$–Energy plane  [^1] By convention energy is denoted by frequency $\omega$ in the expression of the neutron structure factor., following a path through reciprocal space traversing a set of high-symmetry points.
+Note that although the points labeled as X, Y, and Z; and as M1, M2 and M3 are equivalent by symmetry, there are paths between them that are inequivalent by symmetry because the space group, $P2_{1}3$, is non-centrosymmetric.
+Data are presented centered at $\mathbf{G}=(0,0,4)$ in fig.
+2 and $\mathbf{G}=(-2,-2,2)$ in fig.
+3 for both $T=100$ K and $T=400$ K, and at $\mathbf{G}=(3,-3,0)$ in fig.
+4 for $T=2$ K and $T=300$ K.
+The DFT-calculated dispersion relation with $a=9.215$, with the energy scale reduced by 5%, is overplotted.
+The calculated structure factor is also shown in panel (c) of all three figures.
+This set of figures illustrate many of the salient points of this work.
+It is clear that the dispersion and cross-section calculated using DFT provide an extremely good level of agreement with the neutron measurements, with significant overlap at most wavevectors, indicating that both the eigenvalues and eigenvectors from the DFT are accurate.
 At most wavevectors a change in temperature of $\sim 300$ K results in very little change in the measured dispersion and $S(\mathbf{Q},\omega)$.
@@
 6.
-Figure 5: Cut at the mid-point of the Y–R trajectory of fig.
-4(a) and (b), i.e.
-at $\mathbf{Q}=(3.25,-2.5,-0.25)$, showing a significant hardening of the first phonon peak from $\sim 3.5$ meV to $\sim 4$ meV.
+Figure 5: Cut at the mid-point of the Y–R trajectory of fig. 4(a) and (b), i.e. at $\mathbf{Q}=(3.25,-2.5,-0.25)$, showing hardening of the first phonon peak from $\sim 3.5$ meV to $\sim 4$ meV.
 Figure 6: Cut at the mid-point of the Z–M2 trajectory of fig.
@@
 Despite the overall excellent agreement between the DFT calculations and the INS measurements shown in figs.
-2, 3 and 4, there are a small number of wavevectors for which there is some disagreement.
-These can be seen by comparing the calculated dispersion and cross-section shown in panel (c) with the data in panels (a) and (b) of all three figures.
+2, 3 and 4, there is some disagreement.
+An example can be seen by comparing the calculated dispersion and cross-section shown in panel (c) with the data in panels (a) and (b) of all three figures.
 For example, we can see that for the lowest energy mode on the path M1 – Y (and M2 – Z, which is equivalent by symmetry) the calculated energy is rather higher than the measurement at 100 K, whereas the other low energy ($<4$ meV) modes along other paths agree rather closely at this temperature.
+Figure 7: Cuts along the energy axis at four high symmetry positions in reciprocal space, centered on different Brillouin zones in order to maximize the cross-section.
+Panel (a) shows a cut at G, panel (b) a cut at Z, panel (c) a cut at M2, and panel (d) a cut at R.
+Data at lower temperatures (2 K in panel (a), 100 K in panels (b)–(d)) are shown in blue whereas data at higher temperatures (300 K in panel (a), 400 K in panels (b)–(d)) are shown in red.
+For all cuts an estimated constant background has been subtracted, followed by correction for the Bose-Einstein population factor in order to allow easier comparison of data taken at different temperatures.
+Solid lines correspond to fits to Gaussians (the number of which is set by the number expected from the DFT calculations), to indicate the peak positions.
 To allow a more quantitative analysis we show constant $\mathbf{Q}$ cuts along the energy axis at various high symmetry wavevectors, G, X, M and R, at different temperatures in fig.
@@
 In all cases the integration range along the orthogonal $q$-axes is 0.04 reciprocal lattice units.
-This finite integration range, necessary when plotting data from time-of-flight INS experiments, means that for modes that are highly dispersive near the wavevector of interest the width of peaks may be broader than the nominal instrumental energy resolution.
+This finite integration range, necessary when plotting data from time-of-flight INS experiments, obviously means that for modes that are highly dispersive near the wavevector of interest the width of peaks may be broader than the nominal instrumental energy resolution.
 This is most apparent for the single peak in the cut at R – referring to fig.
 3 one can see that the dispersion is very steep until quite close to the R point.
-Figure 7: Cuts along the energy axis at four high symmetry positions in reciprocal space, centered on different Brillouin zones in order to maximize the cross-section.
-Panel (a) shows a cut at G, panel (b) a cut at Z, panel (c) a cut at M2, and panel (d) a cut at R.
-Data at lower temperatures (2 K in panel (a), 100 K in panels (b)–(d)) are shown in blue whereas data at higher temperatures (300 K in panel (a), 400 K in panels (b)–(d)) are shown in red.
-For all cuts an estimated constant background has been subtracted, followed by correction for the Bose-Einstein population factor in order to allow easier direct comparison of data taken at different temperatures.
-Solid lines correspond to fits to Gaussians (the number of which is set by the number expected from the DFT calculations), to indicate the peak position.
-NTE in ZrW2O8 is reflected in the phonon dispersion relation via the mode Grüneisen parameters, $\gamma_{k}$. For a phonon mode $k$ with frequency $\omega_{k}$ this is defined as
-$\gamma_{k}=-\frac{V}{\omega_{k}}\frac{\partial\omega_{k}}{\partial V}$ (1)
-where $V$ is the volume of the unit cell.
-It was by analysis of the mode Gruüneisen parameters obtained from DFT calculations [28, 29, 30, 31] that it was proposed that NTE in ZrW2O8 is dominated by a collection of low energy phonon modes of varying character.
-However, the way mode Grüneisen parameters are determined from DFT is rather different to what actually happens when warming and cooling a sample.
-In the DFT, the lattice dynamics are calculated at zero temperature in the quasi-harmonic approximation for a series of different fixed lattice parameters close to the ground state parameters determined from a minimization of the overall forces and energies.
-This therefore corresponds to the lattice contracting or expanding under pressure / stress, rather than warming / cooling.
-In reality, the warming will result in an increased amplitude of lattice vibrations, which is not accounted for in the quasi-harmonic approximation, with the consequence that anharmonic effects may become more significant at higher temperatures.
-This point has been discussed extensively by Mittal et al [30, 12], who have calculated this effect for a few individual modes at particular wavevectors in ZrW2O8 .
-Some attempt has been made to quantify the effect more broadly using Raman spectroscopy data [17, 18], which has indicated that as well as having large negative Grüneisen parameters, the lowest energy modes also have the largest anharmonicity.
-There are two effects which govern the behavior of a phonon mode energy with temperature.
-On the one hand, in a conventional material, thermal expansion (on increasing temperature) gives rise to longer bonds between atoms which are weaker and hence give rise to lower energy phonon modes.
-On the other hand, as temperature increases the amplitudes of normal modes increase and intrinsic anharmonicity comes into play.
-This would then yield phonon modes that harden slightly with increasing temperature [16].
-However, in practice in conventional materials this effect is often much smaller than the softening due to thermal expansion and so is masked.
-In an NTE material one might imagine that if the lattice expansion effect is opposite to normal, then both of the effects described above would add together and result in phonon mode hardening substantially with increasing temperature.
-Or even in a RUM scenario, in which the bonds between atoms to not change length and hence strength, the intrinsic anharmonicity would be left over to give mode hardening, albeit perhaps smaller in magnitude.
-Neutron measurements of the phonon density of states in ZrW2O8 [24] indeed indicate that most of the low energy modes in ZrW2O8 do indeed harden on warming.
-Examining fig.
-7, that shows cuts along energy at several key wavevectors in different Brillouin zones, we can gain greater insight into this.
-At the G-point (fig.
-7(a)) at 100 K, we observe one optic mode at $5.10\pm 0.03$ meV with a FWHM of $0.41\pm 0.06$ meV.
-On warming to 400 K the mode is fitted to be centered at $5.12\pm 0.05$ meV, with a FWHM of $0.85\pm 0.1$ meV.
-Our calculations, and those of other groups, indicate that one might expect to see two modes here (see fig.
-4, for example).
-However they are close in energy and dispersive so may not be possible to resolve and may be broadened by additional instrumental effects beyond the bare energy resolution (which is $\sim 0.23$ meV at this energy).
-Certainly, the peak is noticeably broadened, most likely due to thermal effects, at 400 K.
-The behavior of this peak contrasts with the expected softening due to the change in lattice alone of -0.2 meV (for both calculated modes), which Gupta et al [30] predict to be compensated such that the expected shifts are $-0.07$ meV and $-0.03$ meV respectively, which is significantly less than the energy resolution of our measurements  [^1] Note that the predictions of Gupta et al [30] refer to a change in temperature of 0 K to 300 K, i.e.
-the same difference in temperature as for our measurements.
-The offset is unlikely to be important, since NTE in ZrW2O8 is linear over this temperature range..
-The Raman measurements of Oishi et al include this mode [18], which they find to be centered at 5.16 meV and largely independent of temperature, consistent with our finding.
-At the Z-point (fig.
-7(b), equivalent to the X- and Y-points) we observe three modes.
-At 100 K they are centered at energies of $3.50\pm 0.05$ meV, $4.43\pm 0.05$ meV and $5.28\pm 0.04$ meV.
-All three were found to have the same FWHM (within error) of $0.49\pm 0.05$ meV.
-At 400 K these peaks are centered at energies of $3.70\pm 0.10$ meV, $4.48\pm 0.09$ meV and $5.44\pm 0.13$ meV respectively.
-The FWHMs were found to be $0.53\pm 0.11$ meV.
-Gupta et al presented predictions for the first two of these modes, with a shift due to the change in lattice of -0.14 and -0.16 meV respectively, that is compensated by thermal effects to give expected shifts of 0.22 meV and 0.68 meV.
-The first of these predictions is consistent with our measurement, whereas the latter is not – the predicted effect of temperature (referred to as explicit anharmonicity) is larger than what is actually measured.
-At the M-point (fig.
-7(c), actually M2 in our notation) we observe three modes.
-At 100 K they are centered at energies of $3.77\pm 0.04$ meV, $4.35\pm 0.05$ meV and $5.60\pm 0.03$ meV.
-All three were found to have the same FWHM (within error) of $0.42\pm 0.04$ meV.
-At 400 K these peaks are centered at energies of $3.58\pm 0.07$ meV, $4.05\pm 0.06$ meV and $5.62\pm 0.04$ meV respectively.
-The FWHMs were found to be $0.34\pm 0.06$ meV.
-For the first two of these modes, Gupta et al predict shifts due to a change in lattice of -0.36 and -0.37 meV respectively, compensated by explicit anharmonicity to give overall predicted shifts of 0.06 and 0.18 meV.
-In contrast, our fits would seem to indicate that both modes soften slightly.
-However, inspecting fig.
-3 we see that care should be taken here – it is not obvious from the colormap that these two modes soften as much as the cuts and fits seem to suggest.
-In fact, what appears to happen is the lowest energy mode which disperses weakly along the Z – M2 trajectory with an energy of $\sim 3.5$ meV seems to steepen slightly on warming, with the effect of bringing some of its spectral weight into the integration region of the cut at 400 K shown in fig.
-7(c).
-Therefore, in the absence of a calculation for the shifts in energy at all $\mathbf{Q}$, we cannot quantify with much accuracy how well the measurement and prediction agree at this wavevector.
-The predictions suggest small shifts at about the level of our experimental resolution, which looks qualitatively reasonable.
-At the R-point (fig.
-7(d)) we observe one mode.
-At 100 K it is centered at an energy of $4.45\pm 0.02$ meV, with a FWHM of $0.56\pm 0.03$ meV.
-At 400 K the peak is centered at an energy of $4.43\pm 0.05$ meV with a FWHM of $0.91\pm 0.11$ meV.
-Here, Gupta et al predict a shift due to the change in lattice of -0.39 meV, which is complemented by a significant explicit anharmonicity of the opposite sign to those at other wavevectors, to give an overall predicted shift of -0.77 meV.
-The large FWHM observed at both temperatures is due to the steep dispersion of the modes near the R-point and the consequent inclusion in the cut integration region of spectral weight away from the R-point itself.
-Notwithstanding this, it is clear from both fig.
-7(d) and fig.
-3 that the mode near R does not soften as much as predicted.
 III.2 DFT calculations
-The DFT calculations were performed at a lattice parameter of 9.215Å, and we applied a global re-scaling of the calculated phonon energies of 95% to give best agreement with the INS measurements  [^2] This was done by doing a least squares fit of the neutron spectra with the DFT-calculated spectra, varying only the scale factor to give the best agreement.
-Coincidentally, the relative difference in lattice parameter between that used for the DFT and that observed experimentally at low temperature is approximately the same as the experimentally observed negative thermal expansion between 100 K and 400 K [9].
+The DFT calculations plotted in figures.
+24 were performed at a lattice parameter of 9.215Å, and we then reduced all calculated phonon energies by 5% to give best agreement with the INS measurements  [^2] This was done by doing a least squares fit of the neutron spectra with the DFT-calculated spectra, varying only the scale factor to give the best agreement.
 We show in fig.
@@
 However, the scaling is not uniform, as can be seen in the inset which shows the trajectory along Z – M2 – Y.
-Here it is clear that the first optic mode (labelled 2 and 4 for the simulation with $a=9.259$  (the optimized lattice parameter in the DFT calculations) and $a=9.215$  respectively) is significantly more dispersive between Z and M2 for the larger lattice parameter.
-This also illustrates why we chose to compare the INS data with the calculations for the unit cell with $a=9.215$  rather than $a=9.259$.
-A global re-scaling of the latter results in a worse agreement with the measurement than the former.
-Figure 8: Comparison of dispersion from DFT calculations with different lattice parameter.
+For instance, the two lowest energy modes for $a=9.210$  and $a=9.215$  are relatively flat in energy, whereas the equivalent modes for $a=9.259$  are much more dispersive.
+Furthermore, one can see that at the M2 point the highest energy mode decreases by $\sim 0.2$ meV on compression of the lattice from $a=9.259$  to $a=9.215$  ($\Delta a=-0.044$) and decreases by even more, $\sim 0.3$ meV, for further compression $\Delta a=-0.005$.
+As noted above, we compared the INS data with the calculations for the unit cell with $a=9.215$  rather than $a=9.259$, which is the zero-stress optimized unit cell dimension in the DFT, or the cell with $a=9.210$.
+Globally rescaling the phonon energies in either of the latter two sets of calculations resulted in worse agreement with the neutron data.
+Figure 8: Comparison of dispersion from DFT calculations with different lattice parameters.
 The reciprocal space path is the same a described in fig.
 2.
-The solid red curve is for calculations with lattice parameter $a=9.259$(zero stress, optimized lattice parameter), and the dashed blue curve is for calculations with $a=9.215$.
+The solid red curve is for calculations with lattice parameter $a=9.259$  (zero stress, optimized lattice parameter), the dotted blue curve is for calculations with $a=9.215$, and the dash-dotted cyan curve is for calculations with $a=9.215$.
 The inset shows a zoomed in section of the dispersion between 3 and 6 meV.
-The modes labelled 1, 2 and 5 correspond to calculations with $a=9.215$, and the modes labelled 3, 4 and 6 correspond to calculations with $a=9.259$.
 The symbols at high symmetry wavevectors correspond to the fitted peak positions for the lower temperature datasets shown in fig.
-7, with diamonds for the G-point, stars for the Z-point, pentagons for the M2-point and a circle at the R-point.
-One might naturally ask if further compression of the lattice in the calculations could yield yet better agreement.
-In fig.
-9 we show calculations with $a=9.215$  and $9.210$.
-Here we see that certain modes, for example on the path Y – R – G, have softened considerably, such that an overall re-scale of energies would once again not yield good agreement with the INS data.
-What this seems to illustrate is that the calculations which agree best with the INS data are those which are on the edge of dynamical stability.
-Indeed, as noted in sec.
-II.1, when the lattice was compressed further in the DFT calculations, a significant number of soft modes appeared, consistent with this interpretation.
-With further compression negative energy modes appear, indicating an eventual change in the ground state structure.
-Figure 9: Comparison of dispersion from DFT calculations with lattice compressed further.
-The reciprocal space path is the same a described in fig.
-2.
-The solid cyan curve is for calculations with lattice parameter $a=9.210$, and the dashed blue curve is for calculations with $a=9.215$.
-Figure 10: The phonon density of states (PDOS) as a function of energy computed using DFT.
-The total PDOS is shown by the black curve, and the element-wise contributions are shown in red (zirconium), green (tungsten) and blue (oxygen).
-Fig.
-10 shows the partial phonon density of states (PDOS) computed from DFT, broken down into the contributions from each of the constituent atomic species.
-The advantage of computing this property is that it is intrinsically averaged across all of reciprocal space and hence offers some insight into the combined effect of many different phonon modes to the material’s thermal properties.
-The disadvantage, as already noted, is that the averaging means that information about particular phonon modes is lost.
-In the energy range probed by this experiment, the PDOS has approximately equal weight from tungsten and oxygen, but very little contribution from zirconium.
-This indicates that the zirconium ions are much more fixed in position, somewhat characteristic of rigid-unit-modes.
-Further decomposition of the PDOS down to the individual atom level (see supplementary information [32]) shows that the contribution from individual tungsten ions is larger than for individual oxygen ions by about a factor two to three at low energies, but the oxygen ions are more numerous so sum up to give a comparable contribution to the PDOS.
-Figure 11: The phonon density of states (PDOS) weighted by the element-specific neutron scattering factors, as a function of energy computed using DFT.
-The total PDOS is shown by the black curve, and the element-wise contributions are shown in red (zirconium), green (tungsten) and blue (oxygen).
-Figure 12: Slices through the $\mathbf{Q}$ energy plane, following the same scheme as fig.
-3.
-The colormaps show the neutron scattering intensity calculated using DFT, decomposed into the contributions from each individual element.
-Panel (a) shows the contribution from zirconium, panel (b) the contribution from tungsten, and panel (c) the contribution from oxygen.
-Note that the intensity scale is the same as that in fig.
-3 and is the same for all three panels.
-Fig.
-11 shows the PDOS with neutron weighting applied (see eq.
-1 in the supplementary information [32]), which allows direct comparison with the neutron structure factor measured in an experiment, such as that in ref.
-24.
-The scattering lengths of all three constituent elements are similar in magnitude, as are their Debye-Waller factors, so the largest effect on the neutron cross-section comes from the division of all the other terms by the atomic mass.
-This gives significantly more weight to the oxygen modes than to tungsten and zirconium, with the consequence that the neutron scattering signal is dominated by oxygen motion.
-We can use a similar procedure for the single crystal plots shown in figs.
-24, i.e.
-of element-wise decomposition of the signal.
-Examples are shown in fig.
-12, for the $(-2,2,2)$ Brillouin zone center, i.e.
-equivalent to fig.
-3.
-As with the PDOS, most of the signal arises from oxygen motions, with contributions from tungsten and zirconium around an order of magnitude smaller.
-However, although the neutron measurements are not directly sensitive to the motion of the heavier metal atoms, the fact that the DFT gives good agreement for the oxygen means that we can be confident that its predictions for the zirconium and tungsten motion are accurate, since each of them are surrounded by a cage of oxygen.
-IV Discussion
-From the results and analysis presented so far, we can see that there is generally very good agreement between the phonon dispersion relations and structure factors computed from DFT, and our inelastic neutron scattering measurements.
-Of course, anharmonicity is essential to describe any kind of thermal expansion, negative or positive.
-We have already cited extensively the predictions made by Gupta et al [30], which were based on a phenomenological theory developed for understanding librational modes at $q=0$ in the $\alpha$ phase of solid nitrogen [49].
-In this approach the phonon modes are decomposed into the displacement and rotation of molecular units (analogous to the RUM approach) and the computed potential rewritten as a polynomial to $4^{\rm{th}}$ order in terms of these coordinates.
-This is clearly an approximation, though it should be acknowledged that it does a reasonable job, estimating correctly that in ZrW2O8 at some wavevectors the effects of implicit and explicit anharmonicity seem to approximately cancel each other.
-It would seem, then, that for the low energy excitations studied here there is a cancellation of two large effects on the energy of the modes.
-Both arise due to anharmonicity, and given the unconventional behavior of ZrW2O8 it is clear that anharmonicity has an important role in governing the material’s thermal properties.
-In order to cancel the mode softening expected from the mode Grüneisen parameters, large amplitude thermal motions are required.
-This would naturally arise in a structure that is especially flexible, with individual atoms subject either only to weak constraints or strongly constrained only in one or two directions.
-Alternative methods exist in addition to DFT of computing the phonon spectra, that include anharmonicity.
-These include the temperature-dependent effective potential (TDEP) method [50, 51, 52] and the self-consistent phonon (SCPH) method [53, 54].
-These would in principle allow us to investigate in detail the effect of temperature and anharmonicity on the phonon modes, and hence NTE.
-However, the calculations to do this for a system like ZrW2O8 would be prohibitively computationally expensive, and are thus outside the scope of the present work.
-Notwithstanding the aforementioned shortcomings of the DFT computations, it should be acknowledged that they perform an admirable job of reproducing many of the details of the INS measurements, both in terms of the eigenvalues (mode energies) and eigenvectors (measured neutron scattering intensities).
-This then gives high confidence that they can be used to give both a qualitative and quantitative commentary on the nature of the atomic displacements involved in the low energy phonon modes measured here.
-This was discussed extensively already by Gupta et al [30], who correlated the translational and rotational motions of quasi-rigid ZrO6 octahedra and WO4 tetrahedra, together with distortions of those polyhedra, with reported distortions found from local structure probes [55].
-We can also have good confidence that the DFT results can be used as a general interpretative tool, e.g.
-in the flexibility analysis of NTE undertaken by Rimmer et al [28].
-Molecular dynamics simulations [29] suggest that the oxygens that bridge Zr–O–W bonds are subject to large transverse vibrations.
-The distortions we observe from our DFT computations indicate that some modes do indeed include such motions.
-Because of the relative rigidity of the WO4 tetrahedra (it is well-known that the W–O bonds are shorter than the Zr–O bonds, so the WO4 tetrahedra are more tightly bonded and stiffer than the ZrO6 octahedra [56]) a consequence of such vibrations is that the lone unbonded oxygen on each tetrahedron is also subject to large vibrations [25].
-Indeed, displacements of the lone oxygen would have a lower energy cost due to weaker bonding.
-The flexibility analysis performed by Rimmer et al [28], using DFT results similar to those reported here, showed that the low energy phonon modes responsible for most of the NTE (those with negative Grüneisen parameters) had largely fixed Zr–O and W–O bond lengths.
-These results together suggest that the tension effect description of NTE offers an appropriate description of the behavior of ZrW2O8 : rigid Zr–O and W–O bonds together with a large transverse vibration of oxygens that bridge Zr and W lead naturally to a reduced distance between Zr and W and hence NTE.
-By validating the DFT that underpins this interpretation against INS data, we are able to give greater credence to it.
-We also note that in the high pressure orthorhombic phase of $\gamma$-ZrW2O8 [57] the lone oxygen becomes bonded to a tungsten ion, making its motion more constrained than in the ambient pressure cubic $\alpha$-ZrW2O8 phase, which is then manifest in a significantly smaller (negative) coefficient of thermal expansion in the high pressure polymorph.
-V Conclusion
-In conclusion, we present a comprehensive set of INS measurements of the phonons in ZrW2O8 , and compare them to predictions from DFT.
-We find that we have to compress the lattice compared to the equilibrium in the DFT in order to get agreement with the neutron data, but that after doing so the agreement is excellent.
-This gives confidence that the overall picture of atomic motions predicted by DFT closely matches the experimental situation, so that DFT can be confidently used as a basis for additional analysis of NTE.
-The calculated mode Grüneisen parameters suggest that most low energy phonon modes should soften on warming as the lattice shrinks, but instead most modes either remain unchanged or harden slightly, as well as broadening somewhat.
-This can be understood in terms of anharmonicity arising from an increased amplitude of atomic motions associated with the phonon modes.
+7, with diamonds for the Gamma - point, stars for the Z - point, pentagons for the M2 - point and a circle at the R - point.
+As noted, a small compression of the lattice from $a=9.215$  to $a=9.210$  yields a rather large change in mode energies, suggesting that around these lattice parameters the calculations are on the edge of dynamical stability.
+Indeed, when the lattice was compressed further in the DFT calculations (not shown), a significant number of soft modes appeared, consistent with this interpretation.
+With further compression negative energy modes appear, indicating a change in the ground state structure.
+Given this non-linear behavior, to make a quantitative prediction for the effect on the phonon energies of temperature-induced lattice compression in this quasi-harmonic approximation, it makes most sense to compare the calculations with $a=9.259$  and $a=9.215$, rather than those with $a=9.215$  and $a=9.210$.
+The relative difference in the former two lattice parameters is 0.48%, which is quite close to the experimentally observed change in lattice parameter on warming from 100 K to 400 K (0.37%).
+Thus, one can take the changes in the phonon dispersion between these two calculations as representative of the changes in the phonon dispersion over the experimentally probed temperature range, based on the Grüneisen model alone.
+The results of this procedure at various high symmetry wavevectors are presented in table 1.
+Table 1: Comparison of the position of the lowest energy peak in the measured data at 100 K and 400 K, and the expected change in peak position based on DFT alone (i.e.
+quasi-harmonic approximation), and the computed change in peak position from the work of Gupta et al [30], decomposed into the contributions from implicit and explicit anharmonicity, from a phenomenological model.
+Wavevector Lowest energy peak at 100 K (meV) Lowest energy peak at 400 K (meV) DFT shift (meV) [this work] Calculated shift [implicit + explicit] (meV) [30]
+G 5.10(3) 5.12(5) -0.55 -0.07 [-0.22 + 0.15]
+Z 3.50(5) 3.70(10) -0.40 0.08 [-0.14 + 0.22]
+M 3.77(4) 3.58(7) -0.78 0.06 [-0.36 + 0.42]
+R 4.45(2) 4.43(5) -0.80 -0.77 [-0.39 - 0.38]
+The DFT calculations presented so far correspond to the lattice contracting or expanding under pressure / stress, rather than warming / cooling.
+In reality, the warming will result in an increased amplitude of atomic vibrations, which is not accounted for in the quasi-harmonic approximation.
+Anharmonic effects may become more significant at higher temperatures.
+This point has been discussed extensively by Mittal et al [30, 12], who have calculated these effects for a few individual modes at particular wavevectors in ZrW2O8 .
+Some attempt has been made to quantify the effect more broadly using Raman spectroscopy data [17, 18], which has indicated that as well as having large negative Grüneisen parameters, the lowest energy modes also have the largest anharmonicity.
+The effects of temperature on phonons can be decomposed into two parts.
+On the one hand a change in lattice parameter with temperature often gives rise to a changed bond length between atoms hence a change in phonon energy.
+This is sometimes referred to as implicit anharmonicity.
+In conventional materials, increased temperature gives longer and weaker bonds, and hence softened phonons.
+On the other hand, as temperature increases phonon-phonon interactions become increasingly important, sometimes also referred to as explicit anharmonicity, and renormalize the phonon energies.
+This often yields phonon modes that harden slightly with increasing temperature [16].
+However, in practice in conventional materials this effect is often much smaller than the softening due to thermal expansion and so is masked.
+In table 1 we compare the positions of peaks in the INS signal at high symmetry wavevectors as a function of temperature, with the expected shift due to a change in lattice parameter alone from the DFT.
+Additionally, we compare these results to phenomenological calculations by Gupta et al [30] that decompose the shift into components arising from implicit and explicit anharmonicity.
+With the exception of the calculation at the R–point, the changes in peak position from the latter are much smaller than those predicted from the DFT alone, and are closer in magnitude to the effects seen in the INS measurements.
+III.3 Machine learned potentials and molecular dynamics calculations
+[…]
+In order to compare the MD results to the INS data, dynamical structure factor calculations were performed with the lattice parameters fixed to the time-averaged values at each temperature.
+Gaussians were then fitted to the peaks in energy scans up to 6 meV at several high symmetry wavevectors to quantify the temperature-dependent shifts in these modes.
+The results are shown in fig.
+9, with the energy scans shown in the appendices figs.
+1417.
+The majority of the modes harden slightly on warming, with the exception of the highest energy one at the M point ($\mathbf{q}=(0.5,0.5,0)$) which softens slightly.
+We note that between 100 K and 400 K the calculated magnitude of mode hardening of $\sim 0.1$ meV is similar in magnitude to what is seen in our INS measurements.
+But between 0 K and 100 K a much greater mode hardening is predicted, notably with the modes at $\mathbf{q}=(0,0.5,0)$ and $\mathbf{q}=(0.25,0.5,0)$ hardening on warming by $>0.5$ meV, which is not observed in the measurements.
+This may, for example, be due to the precise choice of functional used for the DFT training data on which the interatomic potentials were based.
+We note that the DFT calculations presented in sec.
+III.2 had to be fine-tuned for best agreement with the data, whereas the DFT training data for the MD calculations were not adjusted in this way.
+Figure 9: Temperature dependence of phonon modes below 6 meV computed using MD with machine-learned potentials.
+Red circles indicate the modes at $\mathbf{q}=(0.5,0.5,0)$ (M1 in the notation of fig.
+2), green squares the modes at $\mathbf{q}=(0.25,0.5,0)$, and blue circles the modes at $\mathbf{q}=(0,0.5,0)$ (Y in the notation of fig.
+2).
+IV Discussion and Conclusions
+From the results and analysis presented, we can see that there is very good agreement between the phonon dispersion relations and structure factors measured in our inelastic neutron scattering experiments, and those computed from DFT in the zero-temperature quasi-harmonic approximation, albeit with a compressed lattice compared to the optimized value.
+The agreement is good throughout reciprocal space, which is important in the context of the view that there is no single phonon mode responsible for NTE in ZrW2O8, but many.
+This therefore supports the usefulness of the DFT results in interpreting the physics of NTE in ZrW2O8.
+We do not observe any significant change in the neutron structure factors as a function of temperature, aside from some modes broadening on warming, so we can be confident that the phonon eigenvectors do not change significantly with temperature.
+The fundamental nature of the atomic motions thus do not change with temperature, which is consistent with DFT.
+On the other hand, the DFT predicts softening of most phonon modes on compression of the lattice, and with NTE such a compression happens on warming.
+But we observe that the majority of phonons do not soften, indeed some harden by an amount ($\sim 0.1$ meV) that is just discernible with our instrumental resolution.
+This is because DFT’s quasi-harmonic calculations only capture the lattice compression on warming, but do not capture the full anharmonicity, which plays an increasing role with increasing temperature.
+Finite temperature effects can be included in MD simulations, which in our case used potentials generated through machine learning applied to DFT.
+Our MD simulations showed that the low energy modes at a few high-symmetry wavevectors do indeed harden slightly on warming.
+A larger hardening than what is observed was predicted between 0 K and 100 K, but as already noted this may be due to the specific choice of functional for the DFT training data for the generation of the interatomic potentials.
+Given that the DFT calculations give good phonon eigenvectors when compared to experiment, and the MD calculations show the right behavior of the phonon eigenvalues as a function of temperature when compared to experiment, it is instructive to examine the underlying atomic motions that are common to both to help us understand NTE in ZrW2O8 .
+Such analysis has been done separately for the DFT [28] and MD [38] respectively.
+The former analysis compares the DFT to a set of simplified flexibility models, and concludes that NTE arises due to a tension effect involving nearly rigid Zr–O and W–O bonds, with low energy phonons likely involving quasi-rigid small whole-body rotations of the WO4 tetrahedra.
+The analysis of the MD shows that with increasing temperature the Zr–O and W–O bond lengths indeed change very little, with a small positive expansion, but the W–O–Zr bond angle changes significantly, that could also be described as a tension effect.
+Our INS data are consistent with both sets of analysis, and both point towards the same basic mechanism of NTE.
 Acknowledgements.
@@
 Raw data are available via ref.
-[58].
+39 and processed data via ref.
+40.
 We thank the Institut Laue-Langevin for the allocation of beamtime for the experiments reported here.
 Computing resources were provided by the STFC Scientific Computing Department’s SCARF cluster.
+AID gratefully acknowledges funding from the Ada Lovelace Centre (ALC) at the Science and Technology Facilities Council (STFC) and from the Norwegian Research Council under grant ”Neutron Scattering and Atomistic Simulations for a SUPERior Understanding of SUPERionic Conduction” (grant no.
+344062).
+Appendix I: Methods
+A: Inelastic neutron scattering
+ZrW2O8 crystallizes in the $P2_{1}3$ cubic space group.
+The single-crystalline samples for these experiments were grown as described previously [41].
+The sample used was composed of one crystal of mass 2.4 g with a mosaic of $\sim 0.3$ degrees, which is substantially less than the divergence of the incident neutron beam in the inelastic neutron scattering experiments described below, and hence unimportant for the experimental resolution.
+The lattice parameter was measured to be 9.176 Å  at 2 K, 9.159 Å  at 100 K, 9.1389 Å  at 300 K, and 9.125 Å  at 400 K during the alignment procedure for the inelastic scattering measurements described below.
+These values are consistent with those reported in the literature for powder samples [9].
+The time of flight (ToF) neutron scattering experiments were performed on the IN5 spectrometer at the Institut Laue-Langevin [32].
+Data were collected with incident neutron energy $E_{i}=8.24$ meV, with the final chopper spun at 16,000 r.p.m.
+to give a resolution at the elastic line of 0.27 meV (full-width half-maximum, FWHM).
+The sample was aligned so that the $(H,H,0)$ and $(0,0,L)$ reciprocal space directions were in the horizontal plane of the instrument.
+The data were collected by scanning the sample orientation about the vertical axis in 0.5 degree steps over a range of 161 degrees, starting from the $(1,1,0)$-axis perpendicular to the incident neutron beam.
+Data were collected at temperatures of 100 K and 400 K, with these temperatures chosen in order to increase the signal via the Bose-Einstein population factor while maintaining a large temperature difference between measurements, but avoiding the structural phase transition at $\sim 425$ K [9].
+Additional data were collected with the sample aligned with the $(H,0,0)$ and $(0,K,0)$ directions in the horizontal plane of the spectrometer, with the sample orientation scanned over a range of 88 degrees in 1 degree steps.
+These data were collected with $E_{i}=10.43$ meV and a final chopper speed of 16,000 r.p.m.
+giving a FWHM energy resolution at the elastic line of 0.35 meV.
+Data with this configuration were collected at temperatures of 2 K and 300 K.
+The raw data were processed (units conversion, normalization of intensities, etc.) using the Mantid software package [42].
+Further processing of the data, and all subsequent visualization and analysis, including convolution of calculated spectra with the instrumental resolution, were performed using the Horace software package [43].
+Calculation of the inelastic neutron scattering cross-section ($S(\mathbf{Q},\omega)$) from the DFT-calculated phonons was done using the Euphonic software package’s Horace plugin [44, 45].
+B: Density functional theory
+DFT calculations were performed using the plane-wave pseudopotential method as implemented in the CASTEP package [33].
+The Perdew-Burke-Ernzehof (PBE) exchange-correlation functional [46] was used with norm-conserving pseudopotentials of the RRKJ type [47] from the Rappe and Bennett library [48], a plane-wave cutoff energy of 750 eV, and a $3\times 3\times 3$ Monkhorst-Pack [49] grid for the Brilloun-Zone of integration of electronic states.
+As also observed by Rimmer et al [28], the equilibrium lattice parameter at zero pressure is 9.259 Å, slightly larger than the measured low temperature lattice constant of $\sim 9.18$, typical of the PBE functional [50].
+The phonon dispersion was calculated using Density Functional Perturbation Theory (DFPT) [51], with a $3\times 3\times 3$ Monkhorst-Pack grid of phonon wavevectors and Fourier interpolation of the resulting dynamical matrices, in order to sample the entire Brillouin zone.
+Additional calculations were performed for fixed unit cells of smaller dimension (9.23, 9.22, 9.215, 9.21 and 9.16 Å), at each of which the structure was relaxed and phonons calculated in the same way as for the fully relaxed structure.
+The calculated stress for each of these cell dimensions was 0.81, 1.13, 1.29, 1.45 and 3.20 GPa respectively.
+In the first four of these cases no negative energy phonon modes were observed, indicating that the computed structure remained dynamically stable.
+For the calculation with a lattice parameter of 9.16 Å a large number of negative energy modes were observed, indicating that for such a compressed lattice the structure becomes unstable.
+From this point onward when we refer to these DFT calculations we mean those with $9.21\leq a\leq 9.23$.
+When comparing the computed dispersion with that measured experimentally, an empirical scale factor was applied to the eigenvalues (energies) but not the eigenvectors (structure factors).
+This was determined by finding the simulation that agrees most closely with the data after such a uniform re-scale: the structure factors from the simulations with different lattice parameters were fitted to the neutron data with a single adjustable parameter (the energy scale).
+For ZrW2O8 the best overall agreement was found for the case of $a=9.215$   and a scale factor $\sim 0.95$.
+This approach was used because changing the lattice parameter for the simulation does not affect the eigenvalues uniformly.
+Additional calculations were performed with the same DFT parameters and a lattice parameter of 9.215 Å, but using the PBEsol functional [50], which has been reported as yielding values for the bulk properties (elastic constants) of ZrW2O8  closer to the experimentally determined values than other functionals [52].
+C: Machine learned potentials and molecular dynamics calculations
+Calculations were performed using the automated potential development (APD) workflow [34], which is open-source and available from gitlab [53].
+APD automates all DFT calculations, potential optimizations and potential simulations, as well as any subsequent cycles of active-learning that may be required to refine the potential.
+In addition APD also generates various properties at the DFT and potential level.
+Note that the DFT calculations performed by APD are distinct from those used elsewhere.
+CASTEP with optimized norm-conserving pseudo-potentials [47, 54] was used as the DFT engine for APD.
+Calculations were performed using two generalized gradient approximation exchange-correlation functionals, both PBE [55] and the PBEsol functional [56].
+For all except the phonon calculations a Monkhorst–Pack grid of 2$\times$2$\times$2 (with 44 atom cell) was used.
+For configuration generation, a relatively low cutoff of 520 eV was used.
+Molecular dynamics (MD) simulations were run at 333 K, 666 K and 1000 K, each for 2000 configurations with a timestep of 7 fs.
+Uncorrelated snapshots (separated by $\sim$ 15 configurations) were then recalculated using a cutoff of 1000 eV, ensuring an accuracy of ±$\pm 1$ meV/atom in energies.
+This cutoff, and the k-points grid, were automatically optimized by APD to achieve this accuracy.
+Moment tensor potentials (MTPs) [57] were then optimized using the MLIP2 software package [58], and fitted to energies, forces and stresses (EFS) from the DFT simulations.
+Five distinct potentials (MTP1-5) were independently optimized.
+Once optimized, MTP1-5 were used to drive LAMMPS (Large-scale Atomic/Molecular Massively Parallel Simulator) molecular dynamics simulations [35].
+LAMMPS jobs were performed up to 1000 K in 100 K increments, each for 10 ps using the Nosé-Hoover thermostat and barostat.
+From these simulations, configurations for which EFS predictions are unreliable were identified using MLIP’s MaxVol algorithm, using an extrapolation grade of 3 as a threshold.
+These configurations were then recalculated using DFT, and subsequently used to reoptimize potentials.
+The procedure was repeated until no configurations exceed an extrapolation grade of 3 during the LAMMPS simulations conducted at all temperatures.
+This is handled separately (and automatically) for each potential, so that each undergoes its own active-learning cycle.
+As an additional post-processing step, dynamic structure factors were computed for a selection of q-points and across a range of temperatures.
+Dynamical structure factors were calculated using the Dynasor open-source framework [59].
+Calculations were performed at high-symmetry points $Y=(0,0.5,0)$, $M=(0.5,0.5,0)$, and the midpoint between these points.
+A total of 200 runs, each lasting 1 ns, were performed for each temperature 100 K, 200 K, 300 K and 400 K.
+Each run included a 200 ps equilibration with the Number Volume Temperature (NVT) ensemble, followed by a 60 ps equilibration and a 400 ps production run with the Number Temperature Energy (NTE) ensemble.
+Structure factors were calculated from each run and subsequently averaged.
+Appendix II: Distinguishing phonon modes via their structure factors
+The coherent inelastic neutron scattering cross-section from single-phonon excitations (creation) is given by:
+$\left(\frac{d^{2}\sigma}{d\Omega dE_{f}}\right)$ $=$ $\frac{k_{i}}{k_{f}}\frac{\sigma_{\rm{coh}}}{4\pi}\frac{(2\pi)^{3}}{v_{0}}\frac{1}{2M}e^{-2W}\times$ (1)
+$\sum_{s}\sum_{\mathbf{G}}\frac{(\mathbf{Q}\cdot\mathbf{e}_{s})^{2}}{\omega_{s}}\langle n_{s}+1\rangle\times$
+$\delta(\omega-\omega_{s})\delta(\mathbf{Q}-\mathbf{q}-\mathbf{G})$
+where $k_{i}$ and $k_{f}$ are the incident and final neutron wavevectors, $\sigma_{\rm{coh}}$ is the coherent neutron cross-section of the material being studied, $v_{0}$ is the sample volume, $M$ the mass of the sample, $W$ the Debye-Waller factor, $\mathbf{Q}=\mathbf{G}+\mathbf{q}$ the excitation wavevector, composed of the Brillouin zone vector $\mathbf{G}$ and the reduced phonon wavevector $\mathbf{q}$, $\mathbf{e}_{s}$ is the phonon eigenvector for mode $s$ with frequency $\omega_{s}$, and $\langle n_{s}+1\rangle$ is the Bose-Einstein population factor.
+The key term is the one involving the dot product of the wavevector $\mathbf{Q}$ and the phonon eigenvector $\mathbf{e}_{s}$, which implies that the scattering cross-section is largest when $\mathbf{Q}$ and $\mathbf{e}_{s}$ are parallel, and zero when they are perpendicular.
+By measuring with many different orientations of $\mathbf{Q}$ all of the phonons will give rise to a signal in the INS measurement.
+Such an approach permits modes which may be similar in energy but with different eigenvectors to be resolved.
+We note that single crystal inelastic neutron scattering measurements of ZrW2O8 have been reported [60], but with data collected at only a few high symmetry points in a small number of Brillouin zones, and without sufficient energy resolution to discern subtle shifts with a high degree of certainty.
+Inelastic neutron scattering experiments with single crystal samples have been performed on the structurally simple NTE materials ReO3 [61] and Cu2O [62] in the 1970s and 1980s, again covering only a few high symmetry points in a small number of Brillouin zones.
+More recently, as time-of-flight neutron spectrometers have developed, the multi-Brillouin zone measurement technique used in this study has been employed on structurally simple compounds exhibiting anomalous thermal expansion, e.g.
+Cu2O [63] and NaBr [64], permitting a detailed analysis of the phonon spectra.
+Appendix III: Calculated phonon dispersion and density of states
+In figure 10 we compare calculations for the PBE and PBEsol functionals with the same lattice parameter of 9.215 Å.
+It can be seen that using the results from the PBEsol functional result in slightly higher phonon energies for the same lattice parameter.
+As with the comparison between different lattice parameters using the PBE functional, shown in the main text, we note that at certain trajectories and points there are more obvious differences.
+For example, the inset to fig.
+10 shows that the modes at the R point are significantly closer together in energy for the PBEsol functional than they are for PBE.
+Figure 10: Comparison of dispersion from DFT calculations with different choices of functional, with the same lattice parameter of $a=9.215$.
+The reciprocal space path is the same as described in fig.
+2 of the main text.
+The solid black curve is for calculations with the PBEsol functional and the dashed blue curve is for calculations with the PBE functional (the same as the dashed blue curve in fig.
+7 of the main text).
+The inset shows a zoomed in section of the dispersion between 3 and 6 meV.
+We therefore conclude that for analysing the INS data, calculations with either the PBE or the PBEsol functional are valid.
+On this basis we chose not to proceed with further simulations using the PBEsol functional (e.g.
+full geometry and lattice parameter optimization) since there would likely be little to be gained over the PBE calculations already performed.
+Qualitatively, it would seem that the modes which are most sensitive to perturbations to the choice of psuedopotential are also the modes that seem to be most difficult for any of the DFT calculations to reproduce, i.e.
+those modes between about 3 and 5 meV.
+These are also the modes with the largest magnitude mode Grüneisen parameters and hence are likely to be those most sensitive to anharmonic effects.
+Next, we consider the phonon density of states (PDOS).
+In fig.
+11 we show the partial phonon density of states (PDOS) computed from DFT, broken down into the contributions from each of the constituent atomic species.
+The advantage of computing this property is that it is intrinsically averaged across all of reciprocal space and hence offers some insight into the combined effect of many different phonon modes to the material’s thermal properties.
+The disadvantage is that the averaging means that information about particular phonon modes is lost.
+Figure 11: The phonon density of states (PDOS) as a function of energy computed using DFT.
+The total PDOS is shown by the black curve, and the element-wise contributions are shown in red (zirconium), green (tungsten) and blue (oxygen).
+In the energy range probed by our INS experiment, the PDOS has approximately equal weight from tungsten and oxygen, but very little contribution from zirconium.
+This indicates that the zirconium ions are much more fixed in position, somewhat characteristic of rigid-unit-modes.
+Further decomposition of the PDOS down to the individual atom level (not shown) confirms that the contribution from individual tungsten ions is larger than for individual oxygen ions by about a factor two to three at low energies, but the oxygen ions are more numerous so sum up to give a comparable contribution to the PDOS.
+Fig.
+12 shows the PDOS with neutron weighting applied (see eq.
+1), which allows direct comparison with the neutron structure factor measured in an experiment, such as that in ref.
+24.
+The scattering lengths of all three constituent elements are similar in magnitude, as are their Debye-Waller factors, so the largest effect on the neutron cross-section comes from the division of all the other terms by the atomic mass.
+This gives significantly more weight to the oxygen modes than to tungsten and zirconium, with the consequence that the neutron scattering signal is dominated by oxygen motion.
+Figure 12: The phonon density of states (PDOS) weighted by the element-specific neutron scattering factors, as a function of energy computed using DFT.
+The total PDOS is shown by the black curve, and the element-wise contributions are shown in red (zirconium), green (tungsten) and blue (oxygen).
+Figure 13: Slices through the $\mathbf{Q}$ energy plane, following the same scheme as fig.
+2 of the main text.
+The colormaps show the neutron scattering intensity calculated using DFT, decomposed into the contributions from each individual element.
+Panel (a) shows the contribution from zirconium, panel (b) the contribution from tungsten, and panel (c) the contribution from oxygen.
+Note that the intensity scale is the same as that in fig.
+3 of the main text and is the same for all three panels.
+We can decompose the simulations shown in figs.
+2 – 4 of the main text, into the contributions from the different atomic species.
+Examples are shown in fig.
+13, for the $(-2,2,2)$ Brillouin zone center, i.e.
+equivalent to fig.
+3 from the main text.
+Most of the signal arises from oxygen motions, with contributions from tungsten and zirconium around an order of magnitude smaller.
+A similar effect is seen for computation of the phonon density of states.
+However, although the neutron measurements are less directly sensitive to the motion of the heavier metal atoms, the fact that the DFT gives good agreement for the oxygen means that we can be confident that its predictions for the zirconium and tungsten motion are accurate, since each of them are surrounded by a cage of oxygen ions.
+Appendix IV: Structure factors from MD calculations
+Here we show the neutron structure factor $S(\mathbf{q},\omega)$ computed from the MD calculations at several fixed values of $\mathbf{q}$, for temperatures of 100, 200, 300 and 400 K.
+The peaks were fitted using Gaussians, and it is the center positions of these that are plotted in fig.
+10 of the main text.
+As well as the peaks moving to higher energies with increasing temperature, one can also see that, as would be expected, they also broaden.
+Figure 14: Energy scans of the MD-computed neutron structure factor $S(\mathbf{q},\omega)$ at fixed wavevectors $\mathbf{q}=(0,0.5,0)$ (blue circles) , $\mathbf{q}=(0.25,0.5,0)$ (green circles), and $\mathbf{q}=(0.5,0.5,0)$ (red circles).
+Data are successively offset on the y-scale by 2 units.
+Fits of the peaks to a Gaussian lineshape are shown by black lines.
+These data are from the MD simulation at $T=100$ K.
+Figure 15: Energy scans of the MD-computed neutron structure factor $S(\mathbf{q},\omega)$ at fixed wavevectors $\mathbf{q}=(0,0.5,0)$ (blue circles) , $\mathbf{q}=(0.25,0.5,0)$ (green circles), and $\mathbf{q}=(0.5,0.5,0)$ (red circles).
+Data are successively offset on the y-scale by 2 units.
+Fits of the peaks to a Gaussian lineshape are shown by black lines.
+These data are from the MD simulation at $T=200$ K.
+Figure 16: Energy scans of the MD-computed neutron structure factor $S(\mathbf{q},\omega)$ at fixed wavevectors $\mathbf{q}=(0,0.5,0)$ (blue circles) , $\mathbf{q}=(0.25,0.5,0)$ (green circles), and $\mathbf{q}=(0.5,0.5,0)$ (red circles).
+Data are successively offset on the y-scale by 2 units.
+Fits of the peaks to a Gaussian lineshape are shown by black lines.
+These data are from the MD simulation at $T=300$ K.
+Figure 17: Energy scans of the MD-computed neutron structure factor $S(\mathbf{q},\omega)$ at fixed wavevectors $\mathbf{q}=(0,0.5,0)$ (blue circles) , $\mathbf{q}=(0.25,0.5,0)$ (green circles), and $\mathbf{q}=(0.5,0.5,0)$ (red circles).
+Data are successively offset on the y-scale by 2 units.
+Fits of the peaks to a Gaussian lineshape are shown by black lines.
+These data are from the MD simulation at $T=400$ K.